What is the opportunity cost of a car in Japan? What is the opportunity cost of an airplane in Japan? 


1
Expert's answer
2022-01-20T10:05:54-0500

Solution:

The opportunity cost is the benefit that would have been obtained if a different option had been chosen.

The opportunity cost of a car in Japan is "\frac{1}{3}" airplane

= "\frac{16}{48}" = "\frac{1}{3}" airplane

 

The opportunity cost of an airplane in Japan is 3 cars

= "\frac{48}{16}" = 3 cars